OpenStudy (anonymous):

What value is added to both sides of the equation x2 – 6x = 10 in order to solve by completing the square?

OpenStudy (anonymous):

\[x^2-6x=10 ?\]

OpenStudy (anonymous):

yess !

OpenStudy (anonymous):

the answer is decimal, take a litle longer to solve

OpenStudy (anonymous):

\[(x-3)^2=10+3^2\]\[x-3=\pm \sqrt{19}\]\[x=3\pm \sqrt{19}\]

OpenStudy (anonymous):

\[\frac{6}{2}=3,3^2=9\]

OpenStudy (anonymous):

there are a nother way to solve

OpenStudy (anonymous):

\[x^2-3x +9=10+9\]

OpenStudy (mertsj):

Add 9 to both sides to complete the square.

OpenStudy (mertsj):

\[x^2-6x+9=10+9\]

OpenStudy (mertsj):

\[(x-3)^2=19\]

OpenStudy (anonymous):

yes,2 way can solve I like the frist way I solve it short cut and faster

OpenStudy (mertsj):

\[x-3=\pm \sqrt{19}\]

OpenStudy (mertsj):

\[x=3\pm \sqrt{19}\]